OpenStudy (anonymous):

Gabrielle's age is three times Mikhail's age. The sum of their ages is 40 . What is Mikhail's age?

OpenStudy (aum):

Let Gabrielle's age be G and Mikhail's age be M G = 3*M --- (1) G + M = 40 ---- (2) Solve for M. From (2): G = 40 - M Put this in (1): 40 - M = 3M Solve for M.

OpenStudy (anonymous):

@aum is it 37?

OpenStudy (aum):

No. 40 - M = 3M solve for M

OpenStudy (anonymous):

40-M=3M -M -M 40=3 __ ___ 3 3 =13.3

OpenStudy (anonymous):

@aum

OpenStudy (aum):

40 - M = 3M + M +M add: 40 = 4M divide by 4: 10 = M or M = 10